# Moduli spaces and breather dynamics of analytic solutions in chiral magnets without Heisenberg exchange interaction

### v1

@@ Abstract
We investigate the special case of the chiral magnet with vanishing Heisenberg exchange energy, whose axisymmetric Skyrmion solution has previously been found. The dynamical equations of this model look like inviscid fluid flow, and by investigating path-lines of this flow we can construct explicit static and dynamic solutions. We find an infinite-dimensional family of static Skyrmions that are related to the axisymmetric Skyrmion by co-ordinate transformations thus discovering a new moduli space, and further infinite-dimensional families of axisymmetric and non-axisymmetric breather-like supercompactons. We call the maps generating such moduli space “contour-preserving” maps.
@@ 1 Introduction
In the case of magnetic Skyrmions, several exact analytic solutions are known in certain limits of the general theory. The BPS (Bogomol’nyi-Prasad-Sommerfield) or supersymmetric limit possesses many solutions, including the single magnetic Skyrmion [6], but a special potential of the form of the Zeeman energy squared is required: the Zeeman energy (external magnetic field) must match exactly half of the effective anisotropy coupling, and it should also be half of the square of the DMI coupling. This precise limit is related to BPS solutions in a related gauged non-linear sigma model [7, 8]. The single magnetic Skyrmion solution was found away from this limit by Döring and Melcher, where the relation between the the coupling of this potential and DMI coupling is relaxed [9]. Finally, in a recent paper by some of the authors [10], several analytic solutions can be applied to the general theory of magnetic Skyrmions in two limits: the lump limit, where the solution is described mainly by the Heisenberg energy and the opposite limit, where the Heisenberg energy is negligible – the restricted limit. The exact solution by Döringer and Melcher is in fact of the lump type, whereas the restricted analytic solutions depend strongly on the potential in question. The restricted limit of the magnetic Skyrme theory contains only the DMI and a potential and the solutions can be compactons, supercompactons or normal solitons with a tail. Compactons are solitons with compact support, that usually have discontinuous derivatives of their fields at the compacton boundary – but a continuous energy density. Supercompactons on the other hand, are also discontinuous in the fields themselves [10].
One may think that the limit of the theory without the Heisenberg exchange interaction is rather strange. Nevertheless, it quite naturally arises in certain supersymmetric models, where such a term appears as the kinetic energy, where supersymmetrization of the theory “eats” the kinetic term [11, 12, 13]. More precisely, this happens for $\mathcal{N}=2$ supersymmetric extensions, because the solution of the auxiliary field contains the kinetic term with the opposite sign, when substituted back into the Lagrangian [^1] This also happens in $3+1$ dimensions for a Skyrme-like theory [14].. Moreover, exact solutions in a certain limit can be used to approximate solutions in the vicinity of that limit, as for example the solutions in the BPS limit were used to find novel solutions for a range of coupling parameters [15].
@@
In Ref. [10], axisymmetric analytic solutions were found for the restricted magnetic Skyrmion. Due to the limitation on the rotation of the magnetic Skyrmion coming from the DMI, it was conceivable that there would not be a nontrivial moduli space. In this paper, however, we find an infinite-dimensional moduli space. We study the equations of the restricted magnetic Skyrmion from the point of view of fluid paths and find that they are described by perfect circles. It turns out that the circles need not be concentric, since the centre can “move” as the radius changes. The only constraint is that the circles may not intersect each other. We call the symmetry connecting one such solution to another a “contour-preserving” map, and these maps generate the moduli space. We thus propose that the restricted magnetic Skyrmion has a moduli space described by the configurations of a “string” which has zero tension up to a certain maximum extension. Reintroduction of an infinitesimal Heisenberg exchange energy gives the string a tension.
This paper is organized as follows. In Sec. 2 we introduce the restricted model and study the moduli space of the analytic solutions from the point of view of fluid paths. In Sec. 3, we consider the time-dependent counterparts that resemble breather-like Skyrmions. In Sec. 4, we contemplate the extension of the model to three dimensions, but conclude that Hopfions do not exist. We conclude in Sec. 5 with a discussion and outlook. We relegate some supplemental material to three appendices.
@@ 2 Fluid flow dynamics and moduli space
where $k$ is the DMI strength and $V(\bm{n})$ is a generic external potential. We assume that the minimum of the energy is a point and impose the boundary condition that the field must approach this point at “large” distances: $\lim_{|\bm{x}|\to\infty}\bm{n}=\bm{N}$. This effectively point-compactifies the plane to a 2-sphere, giving the topology to the Skyrmions, $\bm{n}:S^{2}\to S^{2}$, which are characterized by the second homotopy group
$\pi_{2}(S^{2})=\mathbb{Z}\ni Q,\qquad Q=\frac{1}{4\pi}\int\mathop{}\!\mathrm{d}^{2}x\;\bm{n}\cdot\partial_{1}\bm{n}\times\partial_{2}\bm{n}\ .$ (2)
Due to the absence of the Heisenberg energy term $E_{2}$, it turns out that the assumption of a solution reaching the boundary condition is not always met. Indeed we will shortly see that there are not only compact solutions, but supercompactons, where the topological charge is fractional (see e.g. appendix. B and Ref. [10]).
@@
where $T_{2}$ represents two-dimensional translations, $\operatorname{O}(2)_{\rm diag}$ includes both $\operatorname{SO}(2)_{\rm diag}$ and the diagonal parity symmetry described above, and $\ltimes$ is the semi-direct product. When the energy involves the Heisenberg energy term $E_{2}$, this is the whole symmetry group. We see below that without it, the full symmetry is in fact much larger.
@@
That means we can switch to a “Lagrangian” description of the Landau-Lifshitz dynamics where we imagine spins being advected by a velocity field $\vec{\nu}=(n_{1},n_{2})$ at speed $2k$, and as we follow a spin, it precesses around the effective magnetic field $\frac{\mathop{}\!\mathrm{d}V}{\mathop{}\!\mathrm{d}\bm{n}}$. Indeed we have that the “convective derivative” is equal to $2k\vec{\nu}\cdot\nabla$, being $\partial_{3}:=0$. Explicitly, we can define a path in (2+1)-dimensional spacetime, $\vec{X}(t)$, such that $\dot{\vec{X}}=2k\vec{\nu}$ and Eq. (4) reduces to
@@
In the fluid description, $\vec{X}(t)$ is the trajectory of an infinitesimal fluid packet, also called a “path-line”. In various cases, we can solve the above equation and explicitly find this path-line and the variation of $\bm{n}$ along it. By putting together path-lines that fill every point in space and time, we can construct dynamical solutions. This is an example of the general method of characteristics [20], which is generally useful in solving first-order nonlinear equations.
For the rest of this paper, we only consider potentials that are symmetric around $\bm{e}_{3}$ (“axisymmetry”), meaning $V(\bm{n})=f(n_{3})$. In the standard cases, the potential corresponds to the combination of Zeeman interaction with strength $h$ and uniaxial anisotropy with strength $u$, i.e.
$f(n_{3})=h(1-n_{3})+u(1-n_{3}^{2})\ .$ (7)
If we define the complex velocity field $\nu:=n_{1}+\mathrm{i}n_{2}$, then we see that the evolution of a spin along the path-line $\vec{X}(t)$ is given by:
@@
$D_{t}\nu=-\mathrm{i}\nu f^{\prime}(n_{3})\ ,$ (9)
so $n_{3}(\vec{X},t)$ is constant and $\nu(\vec{X},t)=-\sqrt{1-n_{3}^{2}}e^{-\mathrm{i}f^{\prime}(n_{3})t}$. We can thus explicitly calculate the curve $\vec{X}$ introducing $X=X_{1}+\mathrm{i}X_{2}$:
$\dot{X}=-2k\sqrt{1-n_{3}^{2}}e^{-\mathrm{i}f^{\prime}(n_{3})t}\implies X(t)=X_{c}+\underbrace{\frac{-\mathrm{i}2k\sqrt{1-n_{3}^{2}}}{f^{\prime}(n_{3})}}_{\mathrm{i}R(n_{3})}e^{-\mathrm{i}f^{\prime}(n_{3})t}\ ,$ (10)
@@
$X(t)=X_{c}+R(n_{3})\big{[}\sin(f^{\prime}(n_{3})t)+\mathrm{i}\cos(f^{\prime}(n_{3})t)\big{]}\ ,$ (11)
so we see that the curve is generically a circle, with radius $R$ depending on $n_{3}$. There are two special cases: if $n_{3}=\pm 1$, then the radius is zero and $\vec{X}(t)$ is a point, $X_{c}$. Notice that in the case of standard potential (7), $f(-1)>f(1)=0$ so the lowest energy state is $n_{3}=1$ everywhere. If on the other hand there are stationary points of $f(n_{3})$ for $-1<n_{3}<1$, then the radius is infinite and $\vec{X}(t)$ is a straight line.
@@ 2.1 Stationary flows
What we have done at this point is to construct a generic path-line for fluid flow. If we first specialize to look for stationary solutions, this path-line is a streamline: all points on this circle instantaneously have the same $n_{3}$, and $\nu$ is such that it is always tangent to the circle. By looking at $\lvert\vec{X}(t)-\vec{X}_{c}\rvert$, the circle will have a signed radius that depends on $n_{3}$, i.e.
$R(n_{3})=\frac{-2k\sqrt{1-n_{3}^{2}}}{f^{\prime}(n_{3})}\ .$ (12)
@@
Figure 1: “Streamlines” of constant $n_{3}$ for positive and negative $R(n_{3})$.
@@ 2.2 Static axisymmetric Skyrmion-like solution
Let us take this a step further and create an axisymmetric Skyrmion-like solution by assembling a series of concentric circles (here we set $X_{c}=0$). For this to be possible, $R(n_{3})$ must be increasing in absolute value as we start from $n_{3}=-1$ (or the circles will intersect). We will see below that in many cases $R(n_{3})$ is not monotonic all the way from $n_{3}=-1$ to $n_{3}=1$, so we must also consider discontinuous solutions where we stop before $n_{3}=1$. In this case, where our constructed solution only reaches a finite radius $R_{\text{max}}$, we limit the domain to be a disc of radius $R_{\text{max}}$, as shown in Fig. 2. Note that if $f^{\prime}(1)\neq 0$, then $R\to 0$ as $n_{3}\to 1$ and it must have reached a maximum at $n_{3}<1$ and we cannot have a whole Skyrmion.
@@
Suppose that $R$ is increasing at least until $R_{\text{Sk}}=R(0)$, the radius canonically defined as the Skyrmion radius, i.e.,
$R_{\text{Sk}}=\frac{2k}{h}\ .$ (14)
In this model, the Skyrmion radius is purely determined by the balance of DMI and Zeeman potential, and the anisotropy only determines the profile around this.
Notice that there exists a special case when $h+2u=0$, then $f^{\prime}(1)=0$ and therefore there is a chance of non-zero $R(1)$. Expanding $n_{3}=1-\frac{\epsilon^{2}}{2}$ and substituting in $f^{\prime}(n_{3})$, we find that $R\to 2k\epsilon/(u\epsilon^{2})$, so the radius is in fact infinite for $\epsilon\to 0$. Let us explicitly calculate this case: $R(n_{3})=\frac{2k\sqrt{1-n_{3}^{2}}}{h(1-n_{3})}=\frac{2k}{h}\sqrt{\frac{1+n_{3}}{1-n_{3}}}$, with $n_{3}=\cos\Theta(r)$ in spherical co-ordinates. We find that
$\Theta(r)=2\arctan\left(\frac{2k}{\sqrt{h}r}\right),$ (15)
@@
Generally, we should first look for the poles of $R$. The condition for a pole is given by $f^{\prime}(n_{3})=0$, i.e.
@@
This implies that in the subset of the polarized phase $h-2\lvert u\rvert>0$, there are no poles within the range $-1<n_{3}<1$, meaning we expect finite $R_{\text{max}}$ to be at $n_{3}<1$, and thus we find supercompactons with fractional topological charge. By contrast, in the canted phase $h+2u<0$ (where $\bm{N}\neq\bm{e}_{3}$) and in the remaining part of the polarized phase $h-2u<0$ we see $R\to\infty$ at $n_{3}=-\frac{h}{2u}$. We then have Skyrmions with fractional topological charge that are not supercompactons in both of these regions of the phase diagram, i.e. they fill the plane and have no discontinuity. To go further, we should look at the maximum of $R(n_{3})$:
@@
$R(n_{3}^{\star})=\frac{2k}{h\sqrt{1-\frac{4u^{2}}{h^{2}}}}\ .$ (18)
@@
The energy of this configuration, according to Derrick scaling, is equal to
$E=-2\pi\int_{0}^{R_{\text{max}}}f(n_{3}(r))r\mathop{}\!\mathrm{d}r\ ,$ (20)
which is always negative, thus indicating a static configuration. Using $R_{\rm max}$ of Eq. (18) and $R$ of Eq. (13), we obtain the analytic result for the energy of the restricted Skyrmion
$E=\frac{k^{2}\pi}{4u^{2}}\left[-\frac{h^{2}-3hu+6u^{2}}{(h-2u)u^{2}}+\frac{(h^{2}-4u^{2})}{u}\left(2\arcsin\left(\frac{2u}{\sqrt{h^{2}-4u^{2}}}\right)+\log\left(1-\frac{4u^{2}}{h^{2}}\right)\right)\right],$ (21)
for $0<u<\frac{h}{2}$ and $h>0$, whereas if $u=0$, the energy simplifies to
$E=-\frac{20\pi k^{2}}{3h}\ ,$ (22)
where the maximal radius now is $R_{\rm max}=2k/h$.
This radially symmetric solution can be also reached by imposing radial symmetry and solving the associated first-order ODE, as shown in appendix B.

### v2

# Moduli spaces and breather dynamics of analytic solutions in Heisenberg exchange-free chiral magnets

@@ Abstract
We investigate the special case of the chiral magnet with vanishing Heisenberg exchange energy, whose axisymmetric Skyrmion solution has previously been found. The dynamical equations of this model look like inviscid fluid flow, and by investigating path lines of this flow we can construct explicit static and dynamic solutions. We find an infinite-dimensional family of static Skyrmions that are related to the axisymmetric Skyrmion by co-ordinate transformations thus discovering a new moduli space, and further infinite-dimensional families of axisymmetric and non-axisymmetric breather-like supercompactons.
@@ 1 Introduction
In the case of magnetic Skyrmions, several exact analytic solutions are known in certain limits of the general theory. The BPS (Bogomol’nyi-Prasad-Sommerfield) or supersymmetric limit possesses many solutions, including the single magnetic Skyrmion [6], but a special potential of the form of the Zeeman energy squared is required: the coupling of the Zeeman energy (external magnetic field) must match exactly half of the effective anisotropy coupling, and it should also be half of the square of the DMI coupling. This precise limit is related to BPS solutions in a related gauged non-linear sigma model [7, 8]. The single magnetic Skyrmion solution was found away from this limit by Döring and Melcher, where the relation between the the coupling of this potential and DMI coupling is relaxed [9]. Finally, in a recent paper by some of the authors [10], several analytic solutions can be applied to the general theory of magnetic Skyrmions in two limits: the lump limit, where the solution is described mainly by the Heisenberg energy and the opposite limit, where the Heisenberg energy is negligible – the restricted limit. The exact solution by Döring and Melcher is in fact of the lump type, whereas the restricted analytic solutions depend strongly on the potential in question. The restricted limit of the magnetic Skyrme theory contains only the DMI and a potential and the solutions can be compactons, supercompactons or normal solitons with a tail. Compactons are solitons with compact support, that usually have discontinuous derivatives of their fields at the compacton boundary – but a continuous energy density. Supercompactons on the other hand, are also discontinuous in the fields themselves [10].
One may think that the limit of the theory without the Heisenberg exchange interaction is rather strange. Nevertheless, it quite naturally arises in certain supersymmetric models, where supersymmetrization of the theory “eats” the kinetic term [11, 12, 13]. More precisely, this happens for $\mathcal{N}=2$ supersymmetric extensions, because the solution of the auxiliary field contains the kinetic term with the opposite sign, when substituted back into the Lagrangian [^1] This also happens in $3+1$ dimensions for a Skyrme-like theory [14].. Moreover, exact solutions in a certain limit can be used to approximate solutions in the vicinity of that limit, as for example the solutions in the BPS limit were used to find novel solutions for a range of coupling parameters [15].
@@
In Ref. [10], axisymmetric analytic solutions were found for the restricted magnetic Skyrmion. Due to the limitation on the rotation of the magnetic Skyrmion coming from the DMI, it was conceivable that there would not be a nontrivial moduli space. In this paper, however, we find an infinite-dimensional moduli space. We study the equations of the restricted magnetic Skyrmion from the point of view of fluid paths and find that they are described by perfect circles. It turns out that the circles need not be concentric, since the centre can “move” as the radius changes. The only constraint is that the circles may not intersect each other. We thus propose that the restricted magnetic Skyrmion has a moduli space described by the configurations of a “string” which has zero tension up to a certain maximum extension. Reintroduction of an infinitesimal Heisenberg exchange energy gives the string a tension.
This paper is organized as follows. In Sec. 2 we introduce the restricted model and study the moduli space of the analytic solutions from the point of view of fluid paths. In Sec. 3, we consider the time-dependent counterparts that resemble breather-like Skyrmions. In Sec. 4, we contemplate the extension of the model to three dimensions, but conclude that Hopfions do not exist. We conclude in Sec. 5 with a discussion and outlook. We relegate some supplemental material to an appendix.
@@ 2 Fluid flow dynamics and moduli space
where $k$ is the DMI strength and $V(\bm{n})$ is a generic external potential. We can impose the boundary condition that the field must approach a minimum $\bm{N}$ of the potential at “large” distances: $\lim_{|\bm{x}|\to\infty}\bm{n}=\bm{N}$. This effectively point-compactifies the plane to a 2-sphere, giving the topology to the Skyrmions, $\bm{n}:S^{2}\to S^{2}$, which are characterized by the second homotopy group $\pi_{2}(S^{2})=\mathbb{Z}$. The element of this group a field $\bm{n}$ represents can be computed from the integral
$Q=\frac{1}{4\pi}\int\mathop{}\!\mathrm{d}^{2}x\;\bm{n}\cdot\partial_{1}\bm{n}\times\partial_{2}\bm{n}\ ,$ (2)
called the ‘topological charge’. However more broadly, we can suppose that if there is a degenerate circle of minima, the field approaches different minima in different directions at infinity, mapping the ‘circle at infinity’ to the circle of minima with a prescribed winding number. This is typically forbidden by the resulting divergence of the Heisenberg energy term, but here it is allowed. The resulting topology is then described by relative homotopy of the map $(D^{2},\partial D^{2})\to(S^{2},S^{1}\xhookrightarrow{}S^{2})$, and the integral above is no longer integer, but still invariant under continuous transformations respecting the boundary conditions. Indeed we will shortly see that there are not only compact solutions, but supercompactons, where the topological charge is fractional (see e.g. Ref. [10]).
@@
where $T_{2}$ represents two-dimensional translations, $\operatorname{O}(2)_{\rm diag}$ includes both $\operatorname{SO}(2)_{\rm diag}$ and the diagonal parity symmetry described above, and $\ltimes$ is the semi-direct product. When the energy involves the Heisenberg energy term $E_{2}$, this is the whole symmetry group.
@@
That means we can switch to a “Lagrangian” description of the Landau-Lifshitz dynamics where we imagine spins being advected by a velocity field $\vec{\nu}=(n_{1},n_{2})$ at speed $2k$, and as we follow a spin, it precesses around the effective magnetic field $\frac{\mathop{}\!\mathrm{d}V}{\mathop{}\!\mathrm{d}\bm{n}}$. Indeed we have that the “convective derivative” is equal to $2k\vec{\nu}\cdot\nabla$, with $\partial_{3}:=0$. Explicitly, we can define a path in (2+1)-dimensional spacetime, $\vec{X}(t)$, such that $\dot{\vec{X}}=2k\vec{\nu}$ and Eq. (4) reduces to
@@
In the fluid description, $\vec{X}(t)$ is the trajectory of an infinitesimal fluid packet, a “path line” in the language of fluid mechanics [20]. In various cases, we can solve the above equation and explicitly find this path line and the variation of $\bm{n}$ along it. By putting together path lines that fill every point in space and time, we can construct dynamical solutions. This is an example of the general method of characteristics [21], which is generally useful in solving first-order nonlinear equations.
For the rest of this paper, we only consider potentials that are symmetric around $\bm{e}_{3}$ (“axisymmetry”), meaning $V(\bm{n})=U(n_{3})$. In the standard cases, the potential corresponds to the combination of Zeeman interaction with strength $h\geq 0$ (without loss of generality) and uniaxial anisotropy with strength $u$, i.e.
$U(n_{3})=h(1-n_{3})+u(1-n_{3}^{2})\ .$ (7)
If we define the complex velocity field $\nu:=n_{1}+\mathrm{i}n_{2}$, then we see that the evolution of a spin along the path line $\vec{X}(t)$ is given by:
@@
$D_{t}\nu=-\mathrm{i}\nu U^{\prime}(n_{3})\ ,$ (9)
so $n_{3}(\vec{X}(t),t)$ is constant along a path line and $\nu(\vec{X}(t),t)=-\sqrt{1-n_{3}^{2}}e^{-\mathrm{i}U^{\prime}(n_{3})t}$. We can thus explicitly calculate the curve $\vec{X}(t)$ introducing $X=X_{1}+\mathrm{i}X_{2}$:
$\dot{X}(t)=-2k\sqrt{1-n_{3}^{2}}e^{-\mathrm{i}U^{\prime}(n_{3})t}\implies X(t)=X_{c}+\underbrace{\frac{-\mathrm{i}2k\sqrt{1-n_{3}^{2}}}{U^{\prime}(n_{3})}}_{\mathrm{i}R(n_{3})}e^{-\mathrm{i}U^{\prime}(n_{3})t}\ ,$ (10)
@@
$X(t)=X_{c}+R(n_{3})\big[\sin(U^{\prime}(n_{3})t)+\mathrm{i}\cos(U^{\prime}(n_{3})t)\big]\ ,$ (11)
so we see that the curve (path line) is generically a circle, with radius $R$ depending on $n_{3}$. There are two special cases: if $n_{3}=\pm 1$, then the radius is zero and $\vec{X}(t)$ is a point, $X_{c}$. Notice that in the case of standard potential (7), U(1)>U(1)=0U(-1)>U(1)=0 so the lowest energy state is $n_{3}=1$ everywhere. If on the other hand there are stationary points of $U(n_{3})$ for $1<n_{3}<1$, then the radius is infinite and $\vec{X}(t)$ is a straight line.
When using the method of characteristics to construct solutions to an equation of motion, discontinuities may occur, and in turn solutions can only satisfy the equation of motion at the integral level, leading to shock conditions. In this paper, we focus on constructing static and dynamical solutions that are continuous in time and space, either for all time or up to a singularity. This means in turn that we must generically restrict the domain to a disc, since the ‘supercompacton’ solutions that are found in this regime are not continuous in the whole plane [10].
@@ 2.1 Stationary flows
What we have done at this point is to construct a generic path line for fluid flow. If we first specialize to look for stationary solutions, this path line is a streamline, defined as an integral curve of the flow at a given instant in time [20]. All points on this circle instantaneously have the same $n_{3}$, and $\nu$ is such that it is always tangent to the circle. By looking at $\lvert\vec{X}(t)-\vec{X}_{c}\rvert$, the circle will have a signed radius that depends on $n_{3}$, i.e.
$R(n_{3})=\frac{-2k\sqrt{1-n_{3}^{2}}}{U^{\prime}(n_{3})}\ .$ (12)
@@
Figure 1: Streamlines of constant $n_{3}$ for positive and negative $R(n_{3})$.
Any static continuous solution to the equation of motion must be assembled from these streamlines in such a way that they do not intersect, and cover the domain. In the case where $R(n_{3})$ has a finite maximum $R_{\rm max}$, this domain must be the disc of radius $R_{\rm max}$, or smaller.
@@ 2.2 Static axisymmetric Skyrmion-like solution
Let us take this a step further and create an axisymmetric Skyrmion-like solution to the static Landau-Lifshitz equation (4) by assembling a series of concentric circles (here we set $X_{c}=0$). For this to be possible, we need $R(n_{3})$ to be a monotonic function (or the circles will intersect). We will see below that in general $R(n_{3})$ is not monotonic all the way from $n_{3}=-1$ to $n_{3}=1$, meaning we can only construct configurations that do not cover the whole target sphere: they have fractional topological charge according to Eq. (2).
@@
Except for the special cases $h=\lvert 2u\rvert$, this function is zero at both 1-1 and $1$ and thus cannot be monotonic. However, we see below that when $h\neq\lvert 2u\rvert$, $R(n_{3})$ is monotonic on two intervals $1,n_{3}^{\star}]$ and $n_{3}^{\star},1]$, with $n_{3}^{\star}$ the location of either a pole or finite maximum of $R$. There are therefore two axisymmetric solutions for a given $(h,u)$, one with $n_{3}=-1$ at the centre and the other with $n_{3}=1$ at the centre, both approaching the same boundary conditions. In case where $n_{3}^{\star}$ is the location of a finite maximum, our constructed solution is a ‘supercompacton’ only reaching a finite radius $R_{\rm max}$, so we limit the domain to be a disc of radius $R_{\rm max}$ to avoid discontinuous solutions, as shown in Fig. 2.
In general, inverting the function (13) gives the profile
$n_{3}(r)=\frac{\rho^{2}\tilde{u}-\sqrt{\left(\rho^{2}+4\right)\tilde{u}^{4}+8\tilde{u}^{2}-\left(\rho^{2}-4\right)}}{\left(\rho^{2}+4\right)\tilde{u}^{2}+4}\ ,\qquad\rho=\frac{r}{k}\sqrt{h^{2}+4u^{2}}\ ,\qquad\tilde{u}=\frac{2u}{h}\ .$ (14)
In the special case $h+2u=0$, $U^{\prime}(1)=0$ and therefore $R(1)$ is not necessarily zero. Expanding $n_{3}=1-\frac{\epsilon^{2}}{2}$ and substituting in $U^{\prime}(n_{3})$, we find that $R\to 2k\epsilon/(u\epsilon^{2})$, so the radius is in fact infinite for $\epsilon\to 0$. Let us explicitly calculate this case: $R(n_{3})=\frac{2k\sqrt{1-n_{3}^{2}}}{h(1-n_{3})}=\frac{2k}{h}\sqrt{\frac{1+n_{3}}{1-n_{3}}}$, with $n_{3}=\cos\Theta(r)$ in spherical co-ordinates. We find that
$\Theta(r)=2\arctan\left(\frac{2k}{hr}\right),$ (15)
@@
Generally, we should first look for the poles of $R$. The condition for a pole is given by $U^{\prime}(n_{3})=0$, i.e.
@@
This implies that in the subset of the polarized phase $h-2\lvert u\rvert>0$, there are no poles within the range $1<n_{3}<1$, meaning we expect finite $R_{\rm max}$ to be at $n_{3}^{\star}<1$, and thus we find supercompactons with fractional topological charge. By contrast, in the canted phase $h+2u<0$ (where $\bm{N}\neq\bm{e}_{3}$) and in the remaining part of the polarized phase h2u<0h-2u<0 we see $R\to\infty$ at $n_{3}=-\frac{h}{2u}$. We then have Skyrmions with fractional topological charge that are not supercompactons in both of these regions of the phase diagram, i.e. they fill the plane and have no discontinuity. To go further, we should look at the maximum of $R(n_{3})$:
@@
$R(n_{3}^{\star})=\frac{2k}{\sqrt{h^{2}-4u^{2}}}\ .$ (18)
@@
We can summarize the two regimes:
$n_{3}^{\star}=\begin{cases}\frac{-h}{2u}&h<\lvert 2u\rvert\,\text{two solutions in whole plane}\\ \frac{-2u}{h}&h>\lvert 2u\rvert\,\text{two solutions on disc}\end{cases}$ (20)
For simplicity, we can refer to these two regimes as ‘infinite $R_{\rm max}$’ and ‘finite $R_{\rm max}$’ respectively.
For the infinite $R_{\rm max}$ solutions, we can ask what the boundary condition at infinity is. When $0\leq h\leq-2u$, $n_{3}^{\star}$ is a minimum of $U(n_{3})$, meaning these solutions approach the degenerate circle of minima as discussed in Sec. 2. On the other hand if $0\leq h\leq 2u$, the field $n_{3}$ tends to the global maximum of the potential at infinity – this calls for discarding the solutions in the parameter region $0\leq h\leq 2u$.
As discussed above, the correct topological description of these solutions is relative homotopy. Here we see that at each given $h,u$ there are two valid static solutions, having the same +1+1 winding around the boundary circle, but one covering the north pole and the other the south pole. Using the formula (2), the topological charge of the configuration with $n_{3}=\pm 1$ at its centre is
$Q^{\pm}=\frac{1}{2}\int_{0}^{\infty}\sin\Theta\partial_{r}\Theta\mathop{}\!\mathrm{d}r=\frac{1}{2}\left(n_{3}(0)-n_{3}^{\star}\right)=\begin{cases}\frac{h}{u}\pm\frac{1}{2}&h<\lvert 2u\rvert\\ \frac{u}{h}\pm\frac{1}{2}&h>\lvert 2u\rvert\end{cases}$. (21)
We see that for our solutions, the two distinct relative homotopies always have different topological charges, although in principle topological charge is a cruder measure.
We expect the energy of the two configurations to be different also, but the DMI term is difficult to calculate explicitly.
These radially symmetric solutions can also be reached by imposing radial symmetry and solving the associated first-order ODE.
From the fluids point of view adopted in this paper, the discontinuous solution in the whole plane found in [10] can be assembled by putting together this static solution on the disc of radius $R_{\rm max}$ together with a uniform solution on the plane minus the disc, noting that since $\vec{\nu}$ is parallel to the boundary of the disc, there is no shock condition.